OpenStudy (therealmeeeee):

What is the five-number summary for the data set? 58 34 47 28 52 60 53 44 36 A. Minimum: 28, Q¹: 34, Q²: 47, Q³: 55.5, Maximum: 60 B. Minimum: 28, Q¹: 35, Q²: 47, Q³: 55.5, Maximum: 60 C. Minimum: 28, Q¹: 35, Q²: 47, Q³: 58, Maximum: 60 D. Minimum: 28, Q¹: 35, Q²: 49.5, Q³: 55.5, Maximum: 60 plzz help

OpenStudy (therealmeeeee):

@tHe_FiZiCx99

OpenStudy (therealmeeeee):

help plzzzzzzzzzzzz

OpenStudy (texaschic101):

put numbers in order 28, 34, 36, 44,47,52,53, 58, 60 minimum -- lowest number -- 28 maximum -- highest number -- 60 median (Q2) = 47 Q1 = (34 + 36) / 2 = 35 Q3 = (53 + 58) / 2 = 55.5 answer is B

OpenStudy (therealmeeeee):

thanks

OpenStudy (texaschic101):

any questions ?

OpenStudy (therealmeeeee):

I might have more

OpenStudy (texaschic101):

ohh....sorry but I am not gonna be on much longer....I have softball practice

OpenStudy (therealmeeeee):

aww k